Commissioning Notes from 1st-2nd September 2008 MST

Observers: J. Hill plus A.Rakich, D.Miller in Tucson, and for the second half J. Storm in Potsdam (our first 3 location observing!)
OSA: A. Ceranski and W. Wack

Summary

Was clear all night, with pretty good seeing at the end. Had a wide variety of telescope wakeup kind of problems. The DSP that does strip encoder interpolation locked up. Dave and Dan got it fixed later in the night. Dave and Dan didn't have much sleep. We measured the new PCS IRTC LEFTZEROPOINT to be -125.9 degrees compared to +19.1 degrees in June (difference=145). We don't understand what accounts for this rotation, but the guider works with this zeropoint. We measured the WFS results to have a difference of 90 degrees compared to June. We don't understand why these two numbers are different. Had a bunch of mysterious problems with PCS and GCS toward the end of the night.

Night time Activities

Times in UT below

Weather finally clears. Only ground fog blowing over the mountain at sunset. Roof humidity sensor reading about 5 points higher than the apparent conditions outside. Roof wind direction isn't managing to cancel out dome rotation.

03:00 Open Dome T=8 degC D=5 degC Tmirror=15 degC (Times in MST below)

Waking up the broken stuff in the telescope

03:30 Preset problems with no polys. Had to restart MCSPU, PCS, MCS. TCSGUI is way cool for this!

03:50 Preset to M5_1469

04:00 Now we are down to not finding a star on the guider. Remember there are big temperature gradients on the telescope.

04:15 Spiral searching without finding star image on IRTC.

04:25 Zglobal=-0.15 (compared to +0.23 where we started) Found star at offset 150, -40.

04:30 Slew to zenith to put cover on entrance to SX mirror cell.

04:48 back to the star, but weird event where PSF gui went red and SX mirror went down (without panic?). Lost ~25 minutes. This happened just at the time when Aaron was ssh'ing into lbtmu102. We made a user panic when the mirror wasn't raising properly. Telescope apparently went down to EL=3 deg for no reason that we understand.

05:00 By now the sky looks to be photometric and D=2 degC while T= 9 degC.

05:10 Try to go back to the star, but elevation axis is doing very strange things.

05:18 Stopall and restart MCSPU.

05:20 Pointo M5_1469 again. First preset fails with bad polys. Didn't find star in spiral search

05:33 Pointto M5_1568

05:40 IIF and irs were somehow hung. Couldn't kill irs, but stopping and restarting IIF freed it.

05:46 Aaron's preset makes a mistake of ~20 degrees in the slew polys.

05:50 Called Dave Ashby. He figures out that the Elevation Strip Encoder Interpolator DSP is hung. Dan Cox is running DSP diagnostics. This is apparently what hit us at 04:48 as well.

06:14 In the meantime, Stephen Hooper has installed a new version of TCS with some new GCS features.

06:18 Everything but ECS is working in the new build.

06:20 pointto M5_1568 again -- Now using only elevation motor encoders rather than strip encoders.

06:25 Stephen has fixed ECS, but it took a long time for the subsystem to start.

06:42 failed to find the star after spiral search

06:43 pointto M5_0835

07:00 found that new PSFL was not successfully communicating with new OSS

07:10 Found star at 40 -30

07:15 Update the IE=-104 CA=-74

07:35 Found a bogus RX value in primary offload in PSFL that was sending garbage to M2. Now M2 collimating correctly.

Measuring the rotator zeropoint

07:38 Tweak IE=-104 CA=-83

07:43 LEFTZEROPOINT from +19.1 to -10.9, then restart PCS. But that was the wrong way, so now +49.1

(LEFTZEROPOINT is in the IRTCBGF section of /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf)

07:53 Not quite enough, so LEFTZEROPOINT=54.1 which is as close as we'll get without serious centroiding (using 20 arcsec offset in RA)

07:58 IE=-100 CA=-94

08:03 Pointto BS9188, but it lands about 15 arcsec from the center -- pointing model doesn't seem very good, but collimation wasn't horrible. Tweak IE=-108 CA=-80.

08:04 setxy -x 0 -y 612.5 -- Doug seeing guider image using "readguidecam", the integrated with GCS command. (Run as telescope on lbtmu107.)

Guide pointto seems to move star wrong way.

08:15 star on acquisition image 310,230 offset 10 0 125,220

Guide pointto fails, it seems to offset the wrong way. acquire000027

08:20 LEFTZEROPOINT=125.9

Guide pointto now guiding, AND displaying on GCSGUI

Testing Active Optics with the GCSGUI

08:23 pressed the start wavefront sensing button on GCSGUI, but it didn't send anything to M1.

08:26 Doug does command line to send to primary (default), but it sends wrong coma to secondary.

08:30 Clear active optics and try again sending to primary (a PSF setting), it sends only one correction

08:34 stop and restart AO loop from GUI to get another correction. Coma runs away and saturates mirror motion.

08:39 zero the gains for Z7, Z8, and start AO again. This time it improved focus. wfsc000010.

08:43 wfsc000012

08:46 Z8 gain back to 1, wfsc000013, 14

08:50 Z7, Z8 gain to -1. Clear active optics. wfsc000015,16 but still runs away.

08:55 Don't send flag=0. Gains to 1. wfsc000017

08:57 Manual +1000 Z7, wfsc000018

08:59 Manual -1000 Z7, wfsc000019

09:00 Manual +1000 Z8, wfsc000020

09:01 Manual -1000 Z8, wfsc000021

Test of Off-axis Guiding

09:08 BS9188 gs=3 at PA=247.96 TelM=GUIDE failed because tertiary mirror reached limit. (Where did that message come from?)

09:11 BS9188 gs=16 at PA=322.4 -- successful guiding off-axis 2.5 arcmin along Y-axis.

09:13 BS9188 gs=16 at PA=270 -- bad polys

09:17 BS9188 gs=16 at PA=10 -- successful guiding off-axis, but on-axis star is off by ~7 arcsec in IRTC.

09:35 Restart MCSPU

stopguiding

AGw Calibration measurements

Collimate

09:46 readGuideCam -e 60000 while Aaron slews rotator from 10 deg to 280 deg. Image is guider000004.fits

09:50 readGuideCam -e 60000 slewing back from 280 deg to 10 deg. guider000005

09:53 readGuideCam -e 60000 with rotator tracking, but no guiding. guider000006, just a bit of image trailing.

10:02 Preset to M103 open cluster PA=180 01:33:16 +60:42:42

10:05 readGuideCam -e 20000 unguided, and uncollimated. 000007

10:07 readGuideCam -e 20000 unguided guider000008

offset 10 0

10:08 readGuideCam -e 20000 unguided guider000009

offset -10 0 (but didn't execute?????)

10:09 readGuideCam -e 20000 unguided guider000010

10:11 readGuideCam -e 20000 unguided guider000011

Preset to PA=90

10:17 readGuideCam -e 20000 unguided guider00012

10:18 stop and start oacserver to include Jesper's new encoder position for the rotator center.

offset 0 0

10:20 readGuideCam -e 20000 unguided guider00013

10:22 readGuideCam -e 600000 slewing from rotator angle 10 to 280 guider000014 (problems with 3 overlapping stars)

10:26 readGuideCam -e 600000 slewing from rotator angle 10 to 280 guider000017

Doug finds 90 deg rotation in AGw wavefronts wrt/M1.

10:37 pointto BS9107 Guide gs=0 PA=0 probe can't move to position

10:39 pointto BS9107 Guide gs=0 PA=180 probe can't move to position, plus rotator problem

home the motor axes

Now agw thinks 0, 612.5 is out of range.

10:46 stop and start oacserver again after Jesper adjusts the encoder calibration.

home the motor axes again, now 0,612.5 is OK

More stuff that we don't completely understand

pointto BS9107 Guide gs=0 PA=180 , but no star in acquisition image, acquire00033

setExpTime 2000

pointto BS9107 Guide gs=0 PA=180, still no star

10:54 pointto BS9108 Guide gs=0 PA=180 no star

11:03 pointto M5_0050 Track

Oops forgot IE, CA when we restarted MCSPU

11:10 Tweak IE=-108 CA=-115

11:13 pointto BS9107 Guide gs=0 PA=180 , star just outside acq field

CA=-100

preset failed because GCS stopped.

11:19 pointto BS9107 Guide gs=0 PA=180 , preset succeed but not guiding?

CA=-108

11:21 pointto BS9107 Guide gs=0 PA=180 , GCS stopped again

11:23 pointto BS9107 Guide gs=0 PA=180 , exposure time too short

stop and restart GCS

setExpTime 2000

11:27 pointto BS9107 Guide gs=0 PA=180 , exptime was back to 0.1 sec

11:29pointto BS9107 Guide gs=0 PA=180 ,

readGuideCam -e 3000

guideimage invalid

11:35 stopAGW startAGW

stop and restart GCS

11:39 readGuideCam -e 6000 takes /tmp/guider000018.fits

11:41 acq 42 is 0.1 sec again

setExpTime 2000

Issue: GCS always takes a 0.1 sec (min)

11:52 IE-118 CA=-113

Now we see the brighter star that was out of the field, and 0.1 sec is OK for acquire.

11:55 GCS stopped when Doug tried to take an IDL WFS image, unknown RPC call.

Guiding fails to lock on the star, the acquisition offset misses the box.

11:57 pointto BS9107 Guide gs=0 PA=180 , succeed in guiding, but Doug kills it again.

Stop and restart GCS with telescope still tracking.

Doug can take a WFS image when GCS isn't guiding.

Preset misses star at edge of acq field

12:03 CA=-120 IE=-122

No star 12:03 CA=-110 IE=-115

12:05 CA=-113 IE=-118

Did spiral search while taking readGuideCam -e 300 25 Saw star go by ^C offset 0 -30 refused because irs says guding not active. Tue Sep 2 05:08:23.178 2008 Offset command failed for left IRTC from IIF

Restarted PCS

12:16 pointto Track

found at 25, 25

CA=-80 IE=-97

12:28 guiding with really excellent images even before we collimate.

Doug program kills GCS again.

12:35 Close Dome in Photometric Skies T=9degC, Wind calm

-- DougMiller - 26 Aug 2008, -- JohnHill - 02 Sep 2008
Topic revision: r5 - 02 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback